# Bruno Ferman  [^2] Sao Paulo School of Economics - FGV    Lucas Finamor [^3] Sao Paulo School of Economics - FGV

### v1

@@ Abstract
Quantitative research relies heavily on coding, and coding errors are relatively common even in published research. In this paper, we examine whether individuals are more or less likely to check their code depending on the results they obtain. We test this hypothesis in a randomized experiment embedded in the recruitment process for research positions at a large international economic organization. In a coding task designed to assess candidates’ programming abilities, we randomize whether participants obtain an expected or unexpected result if they commit a simple coding error. We find that individuals are 20% more likely to detect coding errors when they lead to unexpected results. This asymmetry in error detection depending on the results they generate suggests that coding errors may lead to biased findings in scientific research.
@@ 1 Introduction
In this paper, we test the hypothesis that the probability of detecting a coding error depends on the outcome the error generates. We do so through a randomized experiment embedded in the recruitment process for research positions at a large international economic organization. As part of a coding task designed to assess candidates’ programming ability, we randomize whether a simple coding mistake leads to an expected or unexpected result. The coding mistake is whether individuals take into account that the value 99 in the outcome variable codes missing values. Failing to take into account the missing values leads to wrong results that can be expected or unexpected, depending on the group candidates were randomly allocated. This design allows us to estimate whether individuals are less likely to detect and correct coding errors when the resulting outcome aligns with their expectations. We find that individuals are 20% more likely to detect coding errors when they generate unexpected results. This asymmetry in error detection suggests that coding mistakes may lead not only to an increase in dispersion that is not captured by usual standard errors but also to bias in empirical research.
@@ 2.1 Setting and sample
The data task was the last component of the first screening performed by the recruiters. Completion of the task was encouraged, but it was not a requirement. [^4] Some positions did not require coding, therefore the coding task was not mandatory. At the start of the test, individuals could decide whether to share their data from the test for research purposes. The decision had no impact on how the data task was used for the recruitment process. Pooling the two waves, we have results for 1,065 task takers who agreed to share their data for research purposes.555 In total, 1,230 task takers completed the data task. Among them, 165 (13.4%) opted to not share their data.
@@ 2.2 Data task
We then presented a hypothetical RCT of an intervention that, inspired by the literature, tailored educational content to students’ appropriate level. We explained that 5th-grade teachers in treated schools received materials and training to implement the tailored program, while instruction in control schools remained unchanged. At this stage, participants were asked to report their best guess of the approximate effect of such an RCT on language proficiency, measured 12 months after the start of the intervention. Using a slider, they could select values between –0.30 and 0.30 standard deviations. Consistent with the anchoring provided by existing evidence, more than 86% of task-takers reported priors between 0.08 and 0.16 standard deviations.
@@ 2.3 Identification
Figure 1 presents the observed outcomes for questions Q3 and Q4 for each of these latent types, depending on whether they are in the treated or in the control group. [^9] This classification would not allow individuals to detect the problem in Q3, but not detect it in Q4. Reassuringly, we find that only 2 out of 1065 subjects presented this pattern. Given that, we have two ways of identifying the proportion of CI types.
@@ 2.4 Fairness and ethical concerns
With questions Q3 and Q4, we still have the issue that CI types assigned to the treatment group would realize the 99 issue in Q3—so they would be able to answer correctly Q3 and Q4, whereas those assigned to the control group would only answer correctly Q4. However, candidates are able to edit their answers before submitting them, this eliminates this problem. More specifically, CI types assigned to the control group would only realize that 99 values represent missing values when answering Q4. And then they would be able to edit their answer to Q3 before submitting it for the screening process. The fact that only final answers were used for the screening process, while we rely on initial responses for the experiment, allows the experiment to be ex-post fair while still providing the relevant information for identification of the proportion of participants who only spot the error when it leads to an unexpected result.
@@ 2.5 Descriptive Statistics
Our total sample comprises 1,065 task takers. Among them, 655 (61.5%) completed the data task in the 2024 wave, and 410 (38.5%) in the 2025 wave. Every time a candidate started the task, they would be randomized into the control or treatment group. A total of 558 task takers (52.4%) were randomized into the control group, and 507 (47.6%) were randomized into treatment. In order to correctly identify whether individuals saw the 99 as coding missing values or not, we depend on individuals knowing how to run a regression. Therefore, we define a subset of our sample, the qualified sample, which correctly estimated an OLS regression in question 2, the question preceding the two questions we use in the experiment. A total of 805 (75.6%) candidates are in the qualified sample. Our pre-analysis plan specified at least 800 observations in the qualified sample.
@@
Table 2 shows descriptive statistics of the sample. The first three columns show the overall mean value, mean for the control group, and mean for the treated group for each variable. The next two columns show the estimated difference and p-value for the regression-adjusted balance test of equality of means. The last column shows the number of observations with non-missing information. We can see that 38.6% of the sample are female, 92.8% have a master’s degree or are enrolled in a master’s program, and 85.9% have already taken an Econometrics course. In terms of computational language, 51.4% use Stata, 29.0% R, and 16.1% Python, and the remaining 3.5% use other software. In terms of beliefs about the effects of the hypothetical RCT, the average value is 0.125 standard deviations, in the middle of the interval of the presented papers. Figure A.1 shows the histogram for these priors. The average score in the initial two coding questions was 4.12 out of 6 possible points. Consistent with the randomization protocol, we do not find significant differences between treated and control groups in terms of these covariates. The p-value of a joint test that means of all these variables are the same between these two groups is 0.857. Table A.1 reproduces these results for the qualified sample. [^11] The p-value of a joint test that means of all covariates are the same between treated and control groups is 0.282 for the qualified sample. We do find, however, statistically significant differences in the proportions of individuals with master’s degree or higher and females. As pre-specified in the PAP, we include these (and the other covariates) as controls in our main specifications. Reassuringly, all results remain similar upon the inclusion of covariates. Additionally, Tables 5 and A.5 show results conditioning on these variables.
@@
Prior effect 0.125 0.126 0.124 -0.002 0.474 956
@@
Score Q1-Q2 4.123 4.034 4.222 0.188 0.144 1065
(2.101) (2.174) (2.016) [0.128]
P-value joint test 0.860
@@ 4.1 Main Result
Table 3 presents the main results. In the first column, we present the estimates using the first estimator (presented in Equation 1). In the first panel, we only add the wave fixed effects. We can see that only 7.6% of the control group identifies the 99 in Q3 (the intercept from Equation 1). For the treatment group, the proportion spotting the error is 1.00pp higher, increasing it to 8.6%, although this estimate is not statistically significant. The next two panels add, sequentially, all the control variables: all demographic controls (Panel B) and test measures (computational language used, and initial proficiency scores) in Panel C. The results are all very similar; the point estimate ranges between 0.95pp and 1.02pp, with relatively large standard errors.
@@
Treat $0.0100$ $0.0143$ $0.0138$ $0.0138$
(s.e.) ($0.0194$) ($0.0076$) ($0.0072$) ($0.0072$)
[p-value] [$0.6090$] [$0.0603$] [$0.0498$] [$0.0554$]
{p-value2} {0.30450.3045} {0.03010.0301} {0.02700.0270} {0.02770.0277}
N Obs $805$ $805$ $805$ $805$
Panel B - Controls for wave and demographics
Treat $0.0095$ $0.0154$ $0.0146$ $0.0145$
(s.e.) ($0.0193$) ($0.0079$) ($0.0074$) ($0.0075$)
[p-value] [$0.6237$] [$0.0507$] [$0.0408$] [$0.0530$]
{p-value2} {0.31190.3119} {0.02540.0254} {0.02270.0227} {0.02650.0265}
N Obs $805$ $805$ $805$ $805$
Panel C - Controls for wave, demographics, and test variables
Treat $0.0102$ $0.0142$ $0.0137$ $0.0137$
(s.e.) ($0.0193$) ($0.0078$) ($0.0073$) ($0.0075$)
[p-value] [$0.5966$] [$0.0679$] [$0.0592$] [$0.0687$]
{p-value2} {0.29830.2983} {0.03400.0340} {0.03170.0317} {0.03430.0343}
N Obs $805$ $805$ $805$ $805$
Intercept $0.076$ $0.005$ - -
@@
  • Notes: The table presents the estimates for the four main estimators, each in one column. The three panels differ according to the control variables included in the estimation procedure: only wave indicators (panel A), adding demographics controls (panel B), and adding test variables (panel C). The numbers in parentheses are the robust standard errors. In brackets and curly brackets, the p-values for the bilateral and unilateral tests. For the third column (combined estimators), p-values were obtained using bootstrap with 30,000 replications. For columns 1 and 2 (OLS estimators), we also display the intercept of the regression (control mean and treated mean, respectively). The intercept in Column 1 provides an estimator for the proportion of candidates who always spot the coding errors, regardless of the results they generate. The intercept in Column 2 provides an estimator for the proportion of candidates that spot the errors if they produce conflicting results.
@@
In the second column, we implement the OLS estimator using the data from question 4. The intercept from this regression reveals that only 0.5% of individuals in the treated group spot the error only in this question (that is, those who detect the error when they see the flipped result). In the control group, the proportion that only spots the error in Q4 is higher, yielding a $\hat{\beta}_{2}$ equal to 1.43pp. It is remarkable how close this estimate is to the estimate using only the first question (1.00pp). Moreover, this estimator is considerably more precise than the previous one. This happens because the proportion of treated individuals who only spot in the second question is very close to zero. Even without additional controls, this estimate is marginally significant with a p-value of 0.0603 for a bilateral test (and statistically significant at 5% when we consider a unilateral test, with a p-value of 0.0301). When we add controls in the next panels, the estimate is very stable, ranging from 1.42pp to 1.54pp, with p-values for the bilateral test in the interval 5.0%–6.8%.
@@
Results in the third column combine optimally the two estimators, given their variances and covariances. The estimate without additional controls is 1.38pp. This is an 18.2% increase over the baseline detection probability. Including all controls does not change the result. The estimates are between 1.37 and 1.46 pp with p-values ranging between 4.1% and 5.9%. The results using the GMM estimator, which also combines both sources of identification, are very similar (column 4).
@@
The results so far used the qualified sample, that is, the sample of individuals who know how to run a regression. Table 4 shows our main estimates for two alternative samples. Column 2 shows the results for the sample of all individuals. As expected, the proportion of control individuals spotting the error in the first question is smaller (5.9% compared to 7.6% in the qualified sample). The point estimate is also smaller (0.86pp versus 1.38pp). That is also expected; we cannot identify whether these new individuals correctly spot the 99 or not, because they likely do not know how to run OLS regressions, as they failed question number 2. Indeed, we find that only 16.9% of the unqualified sample provided a correct OLS point estimator (whether taking the 99 values into account or not), compared to 85.5% for the qualified sample. In the third column, we drop from the qualified sample individuals who report negative values for the prior effect, as pre-specified in the PAP. The results look very similar, as only 3 individuals in the sample reported negative values. Lastly, in the fourth column, we restrict the qualified sample to the individuals who had priors close to the presented studies. Here, we consider those who reported priors in the interval 0.08–0.16 standard deviations. The point estimate is slightly larger, 1.53pp versus 1.38pp in the baseline estimation. We would expect this number to be larger as these individuals have priors aligned with the literature, and were, therefore, expecting positive results with the same magnitude that we present them.
@@
Effect $0.0138$ $0.0086$ $0.0138$ $0.0153$
(s.e.) ($0.0072$) ($0.0057$) ($0.0072$) ($0.0081$)
[p-value] [$0.0498$] [$0.1285$] [$0.0510$] [$0.0534$]
{p-value2} {0.02700.0270} {0.06360.0636} {0.02910.0291} {0.02990.0299}
N Obs $805$ $1.065$ $802$ $704$
Spotted First $0.076$ $0.059$ $0.076$ $0.084$
Spotted Second $0.005$ $0.006$ $0.005$ $0.006$
@@ 4.2 Heterogeneity
In this section, we investigate whether we have evidence of heterogeneous effects along some dimensions we observe in the data. For this exercise we use the entire sample for two main reasons. First to have larger groups when we split the analysis in sub-samples. Second, some exercises aim to compare individuals with lower or higher coding ability and skills, and therefore it would be inconsistent to already select on those that scored some screening questions correctly, as we do in the qualified sample. Nevertheless, Table A.4 show the same results for the qualified sample. The first two columns of Table 5 show results separately for men and women. We see very similar results, with women displaying slightly lower baseline detection probability and slightly larger treatment effects.
@@
Gender Clustered ses? Score in Q1-Q2
Men Women No Yes Below median Above median
(1) (2) (3) (4) (5) (6)
Effect $0.0074$ $0.0126$ 0.0041-0.0041 $0.0134$ $0.0045$ $0.0130$
(s.e.) ($0.0063$) ($0.0109$) ($0.0073$) ($0.0071$) ($0.0074$) ($0.0082$)
[p-value] [$0.2191$] [$0.2545$] [$0.4469$] [$0.0545$] [$0.5269$] [$0.0936$]
{p-value2} {0.10620.1062} {0.12570.1257} {0.84660.8466} {0.02830.0283} {0.25190.2519} {0.05140.0514}
N Obs $645$ $405$ $257$ $808$ $533$ $532$
p-value (across subsamples) 0.680 0.087 0.442
Spotted-1st $0.066$ $0.050$ $0.007$ $0.076$ $0.028$ $0.091$
Spotted-2nd $0.003$ $0.010$ $0.008$ $0.005$ $0.008$ $0.004$
  • Notes: Table presents the results for the combined estimator for different samples. Each column represents the estimates for a different sub-sample, displaying heterogeneity by gender (columns 1 and 2), by whether they clustered the standard errors in the screening questions (3 and 4), and below/above the median score in the first two questions (5 and 6). The numbers in parentheses are the robust standard errors. In brackets and curly brackets, the p-values for the bilateral and unilateral tests were obtained using bootstrap with 30,000 replications.

### v2

@@ Abstract
Quantitative research relies heavily on coding, and coding errors are relatively common even in published research. In this paper, we examine whether individuals are more or less likely to check their code depending on the results they obtain. We test this hypothesis in a randomized experiment embedded in the recruitment process for research positions at a large international economic organization. In a coding task designed to assess candidates’ programming abilities, we randomize whether participants obtain an expected or unexpected result if they commit a simple coding error. We find that individuals are almost 20% more likely to detect coding errors when they lead to unexpected results. This asymmetry in error detection depending on the results they generate suggests that coding errors may lead to biased findings in scientific research.
@@ 1 Introduction
In this paper, we test the hypothesis that the probability of detecting a coding error depends on the outcome the error generates. We do so through a randomized experiment embedded in the recruitment process for research positions at a large international economic organization. As part of a coding task designed to assess candidates’ programming abilities, we randomize whether a simple coding mistake leads to an expected or unexpected result. The coding mistake is whether individuals take into account that the value 99 in the outcome variable codes missing values. Failing to take into account the missing values leads to wrong results that can be expected or unexpected, depending on the group candidates were randomly allocated. This design allows us to estimate whether individuals are less likely to detect and correct coding errors when the resulting outcome aligns with their expectations. We find that individuals are almost 20% more likely to detect coding errors when they generate unexpected results. This asymmetry in error detection suggests that coding mistakes may lead not only to an increase in dispersion that is not captured by usual standard errors but also to bias in empirical research.
@@ 2.1 Setting and sample
The data task was the last component of the first screening performed by the recruiters. Completion of the task was encouraged, but it was not a requirement. [^4] Some positions did not require coding, therefore the coding task was not mandatory. At the start of the test, individuals could decide whether to share their data from the test for research purposes. The decision had no impact on how the data task was used for the recruitment process. Pooling the two waves, we have results for 1,036 task takers who started the data task and agreed to share their data for research purposes. [^5] In total, 1,171 task takers started the data task and answered the initial questions. Among them, 135 (11.5%) opted to not share their data.
@@ 2.2 Data task
We then presented a hypothetical RCT of an intervention that, inspired by the literature, tailored educational content to students’ appropriate level. We explained that 5th-grade teachers in treated schools received materials and training to implement the tailored program, while instruction in control schools remained unchanged. At this stage, participants were asked to report their best guess of the approximate effect of such an RCT on language proficiency, measured 12 months after the start of the intervention. Using a slider, they could select values between –0.30 and 0.30 standard deviations. Consistent with the anchoring provided by existing evidence, more than 87% of task-takers reported priors between 0.08 and 0.16 standard deviations.
@@ 2.3 Identification
Figure 1 presents the observed outcomes for questions Q3 and Q4 for each of these latent types, depending on whether they are in the treated or in the control group. [^9] This classification would not allow individuals to detect the problem in Q3, but not detect it in Q4. Reassuringly, we find that only 2 out of 1,036 subjects presented this pattern. Given that, we have two ways of identifying the proportion of CI types.
@@ 2.4 Fairness and ethical concerns
With questions Q3 and Q4, we still have the issue that CI types assigned to the treatment group would realize the 99 issue in Q3—so they would be able to answer correctly Q3 and Q4, whereas those assigned to the control group would only answer correctly Q4. However, candidates are able to edit their answers before submitting them, which eliminates this problem. More specifically, CI types assigned to the control group would only realize that 99 values represent missing values when answering Q4. And then they would be able to edit their answer to Q3 before submitting it for the screening process. The fact that only final answers were used for the screening process, while we rely on initial responses for the experiment, allows the experiment to be ex-post fair while still providing the relevant information for identification of the proportion of participants who only spot the error when it leads to an unexpected result.
@@ 2.5 Descriptive Statistics
Our total sample comprises 1,036 task takers. Among them, 633 (61.1%) completed the data task in the 2024 wave, and 406 (39.9%) in the 2025 wave. Every time a candidate started the task, they would be randomized into the control or treatment group. A total of 527 (50.9%) task takers were randomized into the control group, and 509 (49.1%) were randomized into treatment. In order to correctly identify whether individuals saw the 99 as coding missing values or not, we depend on individuals knowing how to run a regression. Therefore, we define a subset of our sample, the qualified sample, which correctly estimated an OLS regression in question 2, the question preceding the two questions we use in the experiment. A total of 807 (77.9%) candidates are in the qualified sample. Our pre-analysis plan specified at least 800 observations in the qualified sample.
@@
Table 2 shows descriptive statistics of the sample. The first three columns show the overall mean value, mean for the control group, and mean for the treated group for each variable. The next two columns show the estimated difference and p-value for the regression-adjusted balance test of equality of means. The last column shows the number of observations with non-missing information. We can see that 37.7% of the sample are female, 93.7% have a master’s degree or are enrolled in a master’s program, and 86.4% have already taken an Econometrics course. In terms of computational language, 50.8% use Stata, 29.5% R, and 15.9% Python, and the remaining 3.8% use other software. The average score in the initial two coding questions is 4.44 out of 6 possible points. In the recruitment process of the partner institute, they also recorded coding proficiency and scored individuals for their knowledge on impact evaluation. Coding score reflects answers to multiple choice questions on the language of their choice (options were Stata, R, or Python). We assigned each individual the corresponding score from the language they choose to complete our data task. We standardized this variable separately for each software language to have mean zero and unitary standard deviation in the control group. Impact evaluation questions measured individuals familiarity and knowledge on impact evaluation and econometric questions. We also standardized this variable to have mean zero and unitary standard deviation in the control group. In terms of beliefs about the effects of the hypothetical RCT, the average value is 0.126 standard deviations, in the middle of the interval of the presented papers. Figure A.1 shows the histogram for these priors. Consistent with the randomization protocol, we do not find significant differences between treated and control groups in terms of these covariates. The p-value of a joint test that means of all these variables are the same between these two groups is 0.463. Table A.1 reproduces these results for the qualified sample. [^11] The p-value of a joint test that means of all covariates are the same between treated and control groups is 0.606 for the qualified sample. We do find, however, statistically significant differences in the proportions of individuals with master’s degree or higher. As pre-specified in the PAP, we include this (and the other covariates) as controls in our main specifications. Reassuringly, all results remain similar upon the inclusion of covariates. Additionally, Tables 5 and A.4 show results conditioning on these variables.
@@
Score Q1-Q2 4.446 4.438 4.454 0.016 0.891 991
(1.828) (1.850) (1.807) [0.116]
Coding Score 0.055 0.000 0.110 0.110 0.109 886
(1.025) (0.998) (1.049) [0.069]
Impact Evaluation Score 0.032 0.000 0.064 0.064 0.328 949
(1.009) (1.000) (1.018) [0.066]
Prior effect 0.126 0.127 0.125 -0.002 0.466 947
@@
P-value joint test 0.463
@@ 4.1 Main Result
Table 3 presents the main results. In the first column, we present the estimates using the first estimator (presented in Equation 1). In the first panel, we only add the wave fixed effects. We can see that only 7.8% of the control group identifies the 99 in Q3 (the intercept from Equation 1). For the treatment group, the proportion spotting the error is 1.03pp higher, increasing it to 8.8%, although this estimate is not statistically significant. The next three panels add, sequentially, all the control variables: all demographic controls (Panel B), data test measures (computational language used, and initial proficiency scores) in Panel C, and screening variables (coding and impact evaluation scores) in Panel D. The results are all very similar; the point estimate ranges between 0.79pp and 1.11pp, with relatively large standard errors.
@@
Treat $0.0103$ $0.0146$ $0.0141$ $0.0141$
(s.e.) ($0.0199$) ($0.0077$) ($0.0073$) ($0.0073$)
[p-value] [$0.6053$] [$0.0596$] [$0.0514$] [$0.0546$]
{p-value2} {0.30260.3026} {0.02980.0298} {0.02780.0278} {0.02730.0273}
Panel B - Adding demographics controls
Treat $0.0100$ $0.0157$ $0.0150$ $0.0149$
(s.e.) ($0.0197$) ($0.0080$) ($0.0075$) ($0.0077$)
[p-value] [$0.6126$] [$0.0508$] [$0.0400$] [$0.0525$]
{p-value2} {0.30630.3063} {0.02540.0254} {0.02130.0213} {0.02620.0262}
Panel C - Adding data test variables
Treat $0.0111$ $0.0145$ $0.0141$ $0.0141$
(s.e.) ($0.0198$) ($0.0079$) ($0.0075$) ($0.0077$)
[p-value] [$0.5738$] [$0.0674$] [$0.0564$] [$0.0664$]
{p-value2} {0.28690.2869} {0.03370.0337} {0.03080.0308} {0.03320.0332}
Panel D - Adding screening test variables
Treat $0.0079$ $0.0130$ $0.0124$ $0.0124$
(s.e.) ($0.0198$) ($0.0076$) ($0.0072$) ($0.0074$)
[p-value] [$0.6878$] [$0.0864$] [$0.0792$] [$0.0949$]
{p-value2} {0.34390.3439} {0.04320.0432} {0.03800.0380} {0.04750.0475}
N Obs $788$ $788$ $788$ $788$
Intercept $0.078$ $0.005$ - -
@@
  • Notes: The table presents the estimates for the four main estimators, each in one column. The four panels differ according to the control variables included in the estimation procedure: only wave indicators, adding demographics controls, adding data test variables (the language chosen and Score Q1-Q2), and adding screening test variables (coding and impact evaluation scores). Robust standard errors in parentheses. In brackets and curly brackets are the p-values for the bilateral and unilateral tests. For the combined estimator, p-values were obtained using bootstrap with 30,000 replications. The intercept in Column 1 provides an estimator for the proportion of candidates who always spot the coding errors, regardless of the results they generate. The intercept in Column 2 provides an estimator for the proportion of candidates that spot the errors if they produce conflicting results.
@@
In the second column, we implement the OLS estimator using the data from question 4. The intercept from this regression reveals that only 0.5% of individuals in the treated group spot the error only in this question (that is, those who detect the error when they see the flipped result). In the control group, the proportion that only spots the error in Q4 is higher, yielding a $\hat{\beta}_{2}$ equal to 1.46pp. It is remarkable how close this estimate is to the estimate using only the first question (1.03pp). Moreover, this estimator is considerably more precise than the previous one. This happens because the proportion of treated individuals who only spot in the second question is very close to zero. Even without additional controls, this estimate is marginally significant with a p-value of 0.0596 for a bilateral test (and statistically significant at 5% when we consider a unilateral test, with a p-value of 0.0298). When we add controls in the next panels, the estimate is very stable, ranging from 1.30pp to 1.57pp, with p-values for the bilateral test in the interval 5.1%–8.6%.
@@
Results in the third column combine optimally the two estimators, given their variances and covariances. The estimate without additional controls is 1.41pp. This is an 18.1% increase over the baseline detection probability. Including all controls does not change the result. The estimates are between 1.24pp and 1.50pp with p-values ranging between 4.0% and 7.9% for the bilateral test and between 2.1%–3.8% for the unilateral ones. The results using the GMM estimator, which also combines both sources of identification, are very similar (column 4).
@@
The results so far used the qualified sample, that is, the sample of individuals who know how to run a regression. Table 4 shows our main estimates for two alternative samples. Column 2 shows the results for the sample of all individuals. As expected, the proportion of control individuals spotting the error in the first question is smaller (6.8% compared to 7.8% in the qualified sample). The point estimate is also smaller (0.97pp versus 1.41pp). That is also expected; we cannot identify whether these new individuals correctly spot the 99 or not, because they likely do not know how to run OLS regressions, as they failed or did not answer question number 2. Indeed, we find that only 27.6% of the unqualified sample provided a correct OLS point estimator (whether taking the 99 values into account or not), compared to 86.8% for the qualified sample. In the third column, we drop from the qualified sample individuals who report negative values for the prior effect, as pre-specified in the PAP. The results look very similar, as only 3 individuals in the sample reported negative values. Lastly, in the fourth column, we restrict the qualified sample to the individuals who had priors close to the presented studies. Here, we consider those who reported priors in the interval 0.08–0.16 standard deviations. The point estimate is slightly larger, 1.54pp versus 1.41pp in the baseline estimation. We would expect this number to be larger as these individuals have priors aligned with the literature, and were, therefore, expecting positive results with the same magnitude that we present them.
@@
Effect $0.0141$ $0.0097$ $0.0141$ $0.0154$
(s.e.) ($0.0073$) ($0.0065$) ($0.0073$) ($0.0082$)
[p-value] [$0.0514$] [$0.1237$] [$0.0498$] [$0.0541$]
{p-value2} {0.02780.0278} {0.06150.0615} {0.02890.0289} {0.02990.0299}
N Obs $788$ $944$ $785$ $697$
Spotted First $0.078$ $0.068$ $0.078$ $0.085$
Spotted Second $0.005$ $0.007$ $0.005$ $0.006$
@@ 4.2 Heterogeneity
In this section, we investigate whether we have evidence of heterogeneous effects along some dimensions we observe in the data. For this exercise we use the entire sample for two main reasons. First to have larger groups when we split the analysis in sub-samples. Second, some exercises aim to compare individuals with lower or higher coding ability and skills, and therefore it would be inconsistent to already select on those that scored some screening questions correctly, as we do in the qualified sample. Nevertheless, Table A.4 show the same results for the qualified sample. The first two rows of Table 5 show the benchmark results for the entire and qualified samples. As the subsequent panels use the entire sample, the first row should be used as a comparison.
@@
Effect Std Error P-value N Obs Spotted First Diff P-value
Benchmarks
aEntire Sample 0.0097 (0.0065) [0.1237] 944 0.068 -
aQualified Sample 0.0141 (0.0073) [0.0514] 788 0.078 -
1. Clustered SE?
aNo -0.0079 (0.0109) [0.3510] 155 0.014 0.0981
aYes 0.0138 (0.0073) [0.0552] 789 0.077
2. Score in Q1–Q2
aBelow Median 0.0060 (0.0088) [0.4852] 452 0.039 0.5384
aAbove Median 0.0137 (0.0089) [0.0987] 489 0.096
3. Coding Score
aBelow Median -0.0021 (0.0100) [0.8312] 405 0.062 0.3205
aAbove Median 0.0112 (0.0089) [0.2089] 432 0.097
4. Impact Evaluation Score
aBelow Median 0.0043 (0.0107) [0.6709] 423 0.087 0.6365
aAbove Median 0.0107 (0.0083) [0.1982] 454 0.063
5. Master’s degree
aYes 0.0109 (0.0069) [0.1094] 888 0.069 -
6. Econometrics course
aYes 0.0111 (0.0074) [0.1290] 830 0.076 -
7. Gender
aMen 0.0083 (0.0071) [0.1267] 578 0.074 0.6713
aWomen 0.0144 (0.0125) [0.2453] 356 0.058
  • Notes: Table presents the results for the combined estimator for different samples. The columns show respectively the point estimate, standard errors (using bootstrap with 30,000 replications) , p-value, number of observations, proportion of individuals spotting the error in the first question (in the control group), and the p-value testing whether the difference of estimates is different than zero. Each row represents the estimates for a different sub-sample. The first two lines are the benchmark results for the entire and qualified samples. Each panel investigates heterogeneity by whether individuals clustered the standard errors in question 2, score in the first two initial questions, coding score, score on the impact evaluation assessment, whether they have a master’s degree, have taken an econometrics course, and by gender. In panels 5 and 6 we do not display results for those without masters or without econometrics because the sample is very small.